In an ionic compound, the size of the ions affects the internuclear distance (the distance between the centers of adjacent ions), which affects lattice energy (a measure of the attractive force holding those ions together). Based on ion sizes, rank these compounds of their expected lattice energy..
Note: Many sources define lattice energies as negative values. Please rank by magnitude and ignore the sign. |Lattice energy| = absolute value of the lattice energy.
RbCl ,RbBr ,Rbl ,RbF
Answer #1
Concepts and reason

Lattice energy can be defined in two ways:

1.The energy required for breaking an ionic solid to its component atoms, and converting the component atoms into gaseous ions is known as lattice energy.

According to this definition, the reaction is always endothermic, and the lattice energy is always positive.

2.The energy released when the gaseous ions combine to form the ionic solid, is known as lattice energy.

According to this definition, the reaction is always exothermic, and the lattice energy is always negative.

Fundamentals

Lattice energy is also called as lattice enthalpy.

U=Q1Q2r0where,latticeenergy=Uchargeofion1=Q1chargeofion2=Q2internucleardistacebetweentheions=r0\begin{array}{l}\\{\rm{U = }}\frac{{{{\rm{Q}}_{\rm{1}}}{{\rm{Q}}_{\rm{2}}}}}{{{{\rm{r}}_{\rm{0}}}}}\\\\{\rm{where,}}\\\\{\rm{lattice}}\,{\rm{energy}}\,{\rm{ = }}\,{\rm{U}}\\\\{\rm{charge}}\,{\rm{of}}\,{\rm{ion}}\,{\rm{1 = }}{{\rm{Q}}_{\rm{1}}}\\\\{\rm{charge}}\,{\rm{of}}\,{\rm{ion}}\,{\rm{2 = }}{{\rm{Q}}_{\rm{2}}}\\\\{\rm{internuclear distace between the ions = }}{{\rm{r}}_{\rm{0}}}\\\end{array}

The factors affecting the lattice energy are:

1.Ionic charge: As the ionic charge of the ion increases, the lattice energy increases.

2.Ionic radius: As the ionic radius of the ion increases, the lattice energy decreases.

As the size of the ion gets smaller, the ions are packed very close together. Hence, a large amount of energy is required to break the bond between them.

Ionicradiusofflourine=133pmIonicradiusofchlorine=181pmIonicradiusofbromine=196pmIonicradiusofiodine=220pm\begin{array}{l}\\{\rm{Ionic}}\,{\rm{radius}}\,{\rm{of}}\,{\rm{flourine = 133pm}}\\\\{\rm{Ionic}}\,{\rm{radius}}\,{\rm{of}}\,{\rm{chlorine = 181pm}}\\\\{\rm{Ionic}}\,{\rm{radius}}\,{\rm{of}}\,{\rm{bromine = 196pm}}\\\\{\rm{Ionic}}\,{\rm{radius}}\,{\rm{of}}\,{\rm{iodine = 220pm}}\\\end{array}

The decreasing order of lattice energy based on its ion sizes are as follows:

RbF>RbCl>RbBr>RbI{\rm{RbF > }}\,{\rm{RbCl > }}\,{\rm{RbBr > RbI}}

Ans:

The decreasing order of lattice energy is:

RbF>RbCl>RbBr>RbI{\rm{RbF > }}\,{\rm{RbCl > }}\,{\rm{RbBr > RbI}}